# Regularized least squares learning with heavy-tailed noise is minimax optimal

### v2

@@ C.4 Concentration of empirical covariance operators
We have the following classical bound on the estimation error of the empirical covariance operators when Proposition C.1 is applied to the centered random operators $\xi_{i}:=\phi(X_{i})\otimes\phi(X_{i})-\mathcal{C}_{\pi}$ which are bounded by $2\kappa^{2}$ in Hilbert–Schmidt norm under 2.1. Note that the operator norm can be bounded by the Hilbert–Schmidt norm.
Corollary C.5 (Sample error of empirical covariance operator).
@@
We now recall a result by [65, Proposition 1].
Proposition C.6 (Empirical inverse, [65], Proposition 1).
@@
${\mathcal{B}}(n,\delta,\alpha):=1+\log^{2}(2/\delta)\mathopen{}\mathclose{{}\left(\frac{2\kappa^{2}}{n\alpha}+\sqrt{\frac{4\kappa^{2}{\mathcal{N}}(\alpha)}{n\alpha}}}\right)^{2}.$ (C.4)
@@
$\mathopen{}\mathclose{{}\left\|(\widehat{\mathcal{C}}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{-1}(\mathcal{C}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})}\right\|_{\mathcal{L}(\mathcal{H})}\leq 2{\mathcal{B}}(n,\delta,\alpha).$ (C.5)
@@
$\frac{1}{n\alpha}\leq{\mathcal{N}}(\alpha)\;,\qquad 16\max\{1,\kappa^{4}\}\log^{2}(2/\delta)\;\frac{{\mathcal{N}}(\alpha)}{\alpha}\leq n,$ (C.6)
@@
Corollary C.7 (Empirical inverse, worst case scenario).
@@
Using ${\mathcal{N}}(\alpha)\leq c\alpha^{-1}$ and plugging it into (C.4), we obtain
@@
Remark C.8 (Worst case scenario).
@@
Under the eigenvalue decay assumption (EVD), the conditions of Proposition C.6 can be simplified.
Corollary C.9 (Empirical inverse under (EVD)).
@@
where $\tilde{D}$ is defined in Lemma A.1.
@@
By Lemma A.1, under (EVD), there exists a constant $\tilde{D}>0$ such that $\mathcal{N}(\alpha)\leq\tilde{D}\alpha^{-p}$. Inserting this bound into (C.4) leads to
@@
We can use Proposition C.6 to bound the weighted norm of a function in $\mathcal{H}$ by an empirically weighted and regularized norm.
Lemma C.10 (Concentration of weighted regularized norm).
@@
$\mathopen{}\mathclose{{}\left\|\mathcal{C}_{\pi}^{s}f}\right\|_{\mathcal{H}}\leq 2^{s}{\mathcal{B}}(n,\delta,\alpha)^{s}\;\mathopen{}\mathclose{{}\left\|(\widehat{\mathcal{C}}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{s}f}\right\|_{\mathcal{H}}$ (C.7)
@@
$\mathopen{}\mathclose{{}\left\|\mathcal{C}_{\pi}^{s}f}\right\|_{\mathcal{H}}\leq\mathopen{}\mathclose{{}\left\|\mathcal{C}_{\pi}^{s}(\mathcal{C}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{-s}}\right\|_{\mathcal{L}(\mathcal{H})}\mathopen{}\mathclose{{}\left\|(\mathcal{C}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{s}(\widehat{\mathcal{C}}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{-s}}\right\|_{\mathcal{L}(\mathcal{H})}\mathopen{}\mathclose{{}\left\|(\widehat{\mathcal{C}}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{s}f}\right\|_{\mathcal{H}}.$ (C.8)
We bound the terms on the right hand side individually. Applying Lemma D.2 to the first term on the right hand side of (C.8), we have
@@
The second term can be bounded with Proposition C.6. ∎
Appendix D Miscellaneous results
We recall a standard result that addresses the saturation property of Tikhonov regularization, see e.g. [55, Example 4.15].
Lemma D.1.
@@ Appendix D Miscellaneous results
Lemma D.2 (Cordes inequality, [66]).

### v3

@@ D.4 Concentration of empirical covariance operators
We have the following classical bound on the estimation error of the empirical covariance operators when Proposition D.1 is applied to the centered random operators $\xi_{i}:=\phi(X_{i})\otimes\phi(X_{i})-\mathcal{C}_{\pi}$ which are bounded by $2\kappa^{2}$ in Hilbert–Schmidt norm under 2.1. Note that the operator norm can be bounded by the Hilbert–Schmidt norm.
Corollary D.5 (Sample error of empirical covariance operator).
@@
We now recall a result by [guo2017learning, Proposition 1].
Proposition D.6 (Empirical inverse, [guo2017learning], Proposition 1).
@@
${\mathcal{B}}(n,\delta,\alpha):=1+\log^{2}(2/\delta)\mathopen{}\mathclose{{\left(\frac{2\kappa^{2}}{n\alpha}+\sqrt{\frac{4\kappa^{2}{\mathcal{N}}(\alpha)}{n\alpha}}}}\right)^{2}.$ (D.8)
@@
$\mathopen{}\mathclose{{\left\|(\widehat{\mathcal{C}}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{-1}(\mathcal{C}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})}}\right\|_{\mathcal{L}(\mathcal{H})}\leq 2{\mathcal{B}}(n,\delta,\alpha).$ (D.9)
@@
$\frac{1}{n\alpha}\leq{\mathcal{N}}(\alpha)\;,\qquad 16\max\{1,\kappa^{4}\}\log^{2}(2/\delta)\;\frac{{\mathcal{N}}(\alpha)}{\alpha}\leq n,$ (D.10)
@@
Corollary D.7 (Empirical inverse, worst case scenario).
@@
Using ${\mathcal{N}}(\alpha)\leq c\alpha^{-1}$ and plugging it into (D.8), we obtain
@@
Remark D.8 (Worst case scenario).
@@
Under the eigenvalue decay assumption (EVD), the conditions of Proposition D.6 can be simplified.
Corollary D.9 (Empirical inverse under (EVD)).
@@
where $\tilde{D}$ is defined in Lemma B.1.
@@
By Lemma B.1, under (EVD), there exists a constant $\tilde{D}>0$ such that $\mathcal{N}(\alpha)\leq\tilde{D}\alpha^{-p}$. Inserting this bound into (D.8) leads to
@@
We can use Proposition D.6 to bound the weighted norm of a function in $\mathcal{H}$ by an empirically weighted and regularized norm.
Lemma D.10 (Concentration of weighted regularized norm).
@@
$\mathopen{}\mathclose{{\left\|\mathcal{C}_{\pi}^{s}f}}\right\|_{\mathcal{H}}\leq 2^{s}{\mathcal{B}}(n,\delta,\alpha)^{s}\;\mathopen{}\mathclose{{\left\|(\widehat{\mathcal{C}}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{s}f}}\right\|_{\mathcal{H}}$ (D.11)
@@
$\mathopen{}\mathclose{{\left\|\mathcal{C}_{\pi}^{s}f}}\right\|_{\mathcal{H}}\leq\mathopen{}\mathclose{{\left\|\mathcal{C}_{\pi}^{s}(\mathcal{C}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{-s}}}\right\|_{\mathcal{L}(\mathcal{H})}\mathopen{}\mathclose{{\left\|(\mathcal{C}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{s}(\widehat{\mathcal{C}}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{-s}}}\right\|_{\mathcal{L}(\mathcal{H})}\mathopen{}\mathclose{{\left\|(\widehat{\mathcal{C}}_{\pi}+\alpha\operatorname{Id}_{\mathcal{H}})^{s}f}}\right\|_{\mathcal{H}}.$ (D.12)
We bound the terms on the right hand side individually. Applying Lemma E.2 to the first term on the right hand side of (D.12), we have
@@
The second term can be bounded with Proposition D.6. ∎
Appendix E Miscellaneous results
We recall a standard result that addresses the saturation property of Tikhonov regularization, see e.g. [EHN96, Example 4.15].
Lemma E.1.
@@ Appendix E Miscellaneous results
Lemma E.2 (Cordes inequality, [Furuta1989]).
@@
The next result is part of [Rio2017, Proposition 3.5]. It allows us to bound moments between $2$ and $q$. For convenience, we provide a detailed, self-contained proof.
Lemma E.3.
Let $q>2$. For all $\rho\in[2,q]$, $a\in\mathbb{R}_{>0}$, and $x\in\mathbb{R}_{>0}$ we have
$(q-2)x^{\rho}\leq(\rho-2)a^{\rho-q}x^{q}+(q-\rho)a^{\rho-2}x^{2}.$ (E.1)
Proof.
We first show that
$(q-2)\mathopen{}\mathclose{{\left(\frac{x}{a}}}\right)^{\rho-2}\leq(\rho-2)\mathopen{}\mathclose{{\left(\frac{x}{a}}}\right)^{q-2}+(q-\rho).$ (E.2)
For this, define for some $y\in\mathbb{R}_{>0}$ and $t\in\mathbb{R}_{>0}$ the function $f:[0,t]\rightarrow\mathbb{R}$ by $f(s)=\exp(\ln(y)s)$. Note that $f$ is convex (since $f^{\prime}(s)=f(s)\ln(y)$, $f^{\prime\prime}(s)=f(s)\ln(y)^{2}>0$). For $s\in[0,t]$ we then get
$y^{s}=f(s)$ $=f((1-s/t)\cdot 0+s/t\cdot t)$
$\leq(1-s/t)f(0)+s/t\cdot f(t)=(1-s/t)+s/t\cdot y^{t}$
and multiplying both sides by $t$ leads to
$ty^{s}\leq sy^{t}+(t-s).$
Setting t=q2t=q-2, $s=\rho-2$, and $y=x/a$ then shows (E.2). Finally, multiplying both sides of this with $x^{2}/a^{2-\rho}$ establishes the result. ∎
Finally, we require the following basic inequality.
Lemma E.4.
Consider the function $f:(0,1)\to\mathbb{R}_{+}$ given by $f(\alpha)=\log(1/\alpha)^{q}\alpha^{s}$ for some $s>0$ and $q>0$. Then we have
$\sup_{\alpha\in(0,1)}f(\alpha)\leq e^{-q}\mathopen{}\mathclose{{\left(\frac{q}{s}}}\right)^{q}.$
Proof.
We may substitute the transformation $x:=\log(1/\alpha)\in(0,\infty)$ and consider $f(x)=x^{q}\exp(-sx)$ instead. Note that $f$ is nonnegative on $(0,\infty)$ and that $f(0)=\lim_{x\to\infty}f(x)=\lim_{x\to 0_{+}}f(x)=0$. The derivative $f^{\prime}(x)=\exp(-sx)(qx^{q-1}-sx^{q})$ has the unique root $x_{\star}=q/s$ and we hence obtain $f(x_{\star})=\exp(-q)(q/s)^{q}$ as the maximal value. ∎